NASA wants to launch its satellite around an asteroid to make observations. It wants the satellite to be over the same location every 24 hours. If Kepler’s constant, k, is 0.5 hr^2/m^3 , what distance from the center of the asteroid must the satellite be placed?


1
Expert's answer
2021-01-21T14:17:24-0500
"k=\frac{T^2}{r^3}\\0.5=\frac{24^2}{r^3}\\r=10.5\ m"